The following function gives the height, (h) meters, of a batted baseball as a function of the time, t seconds, since the ball was hit.
h=-6(t-2.5)^2+38.5
How many seconds after it was hit did the ball hit the ground, to the nearest second?
------------------------------------
I have made an equation of 0=-6(t-2.5)^2+38.5 and then simplified it out to 0=-6t^2+30t+1 but don't know how to solve for t. i have looked at step by step calculators on a bunch of different sites but all of the instructions involved things that i have not learned yet. The answer in the back of the book is 5 and that's what all of the calculators say but i don't know how to get that answer. I just wanted to know if there is an easier way to solve this equation than the calculators give me. Thanks.

Answer by scott8148(6628)

start with the equation you made

subtracting 38.5 ___ -38.5 = -6(t - 2.5)^2

dividing by -6 ___ 6.417 = (t - 2.5)^2

taking square root ___ ±2.533 = t - 2.5 ___ remember square roots can be positive or negative

adding 2.5 ___ 5.033 = t ___ AND ___ -.033 = t

5 is the answer you are looking for

but the -.033 isn't just junk
___ it represents the time that the ball would have STARTED from the ground